one leg of a right riangle is 1/3m shorter than the other leg. if the hypotenuse of the triangle is 5/3m, find the lengths of both legs.
Answer by checkley71(8403)

x^2+(x-1/3)^2=5/3^2
x^2+x^2-2/3x+1/9=25/9
2x^2-2/3x+1/9-25/9=0
2x^2-2/3x-24/9=0
(2x+2)(x-4/3)=0
x-4/3=0
x=4/3 answer for one of the equal legs.
proof:
4/3^2+(4/3-1/3)^2=5/3^2
16/9+1^2=25/9
16/9+9/9=25/9
25/9=25/9